OpenStudy (anonymous):

Write the equation of the line with an undefined slope, passing through the point (2, 5)

OpenStudy (anonymous):

y = mx + b. Find b by replacing x = 2 and y = 5 in the equation. 5 = 2m + b --> b = 5 - 2m Equation of the line: y = mx + (5 - 2m)